# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 3.5 Vector fields tangent to leaf schemes
The formulation of 3.2 with the module $\Theta_{{\sf T},L}$ arises the question how big this module is? For instance, if $\Theta_{{\sf T},L}=0$ then the hypothesis of 3.2 is automatically satisfied and one might doubt this conjecture. From another side, if ${t}$ is a smooth point of ${\cal F}(\Omega)$ then $L$ is a general leaf and $\Theta\subset\Theta_{{\sf T},L}$. In this section we gather some statements about $\Theta_{{\sf T},L}$.
Let ${\cal F}(\Omega)$ be a foliation in an ${\mathfrak{R}}$-scheme ${\sf T}$, $L$ be a leaf scheme of ${\cal F}(\Omega)$ defined over a larger ring $\check{\mathfrak{R}}$) and through an ${\mathfrak{R}}$-valued point ${t}$ of ${\sf T}$.
@@
(32) $\bar{L}:={\rm Zero}\left(\bar{}{{\cal I}}\right),\ \ \bar{}{{\cal I}}:={\cal O}_{{\sf T}}\cap{{\cal I}}.$
@@
${\cal O}_{{\sf T}}\cap{{\cal I}}=\{0\},$
which means that ${{\cal I}}$ has not algebraic elements.
Note that the concept of Zariski closure depends on the underlying ring $\check{\mathfrak{R}}$. For instance $L$ might be Zariski dense for ${\mathfrak{R}}$ but not for an extension $\check{\mathfrak{R}}$ of ${\mathfrak{R}}$.
@@
Let ${\sf T}$ be an ${\mathfrak{R}}$-scheme and ${t}$ be an ${\mathfrak{R}}$-valued point of ${\sf T}$, not necessarily smooth, and $L$ be a leaf scheme of ${\cal F}(\Omega)$ through ${t}$. We have
(33) $\Theta_{{\sf T},L}=\{{\sf v}\in\Theta_{{\sf T}}\ |\ {\sf v}(\Omega)\subset\bar{}{{\cal I}}\}.$
We may call the right hand side of (33) the dual of $\Omega$ along the Zariski closure of $L$. It is a kind of surprising that we do not need to insert ${\sf v}(\bar{}{{\cal I}})\subset\bar{}{{\cal I}}$ in (33) as it will be clear in the proof.
@@
Proof of $\subset$: If ${\sf v}\in\Theta_{{\sf T},L}$ then ${\sf v}({{\cal I}})\subset{{\cal I}}$, and since ${\sf v}$ is algebraic, that is ${\sf v}({\cal O}_{\sf T})\subset{\cal O}_{\sf T}$), we get ${\sf v}(\bar{}{{\cal I}})\subset\bar{}{{\cal I}}$. Moreover, by definition of a leaf scheme we have $\Omega={\cal O}_{\sf T}d{{\cal I}}$ modulo ${{\cal I}}\Omega_{\sf T}$ (and after tensoring with holomorphic functions). Taking ${\sf v}$ from both sides we have ${\sf v}(\Omega)\subset{{\cal I}}$, and since ${\sf v}(\Omega)\in{\cal O}_{\sf T}$, we get ${\sf v}(\Omega)\subset\bar{}{{\cal I}}$.
Proof of $\supset$: Let ${\sf v}\in\Theta_{{\sf T}}$ with ${\sf v}(\Omega)\subset\bar{}{{\cal I}}$. By definition of a leaf scheme $d{{\cal I}}$ is in $\Omega\otimes_{{\cal O}_{\sf T}}{\cal O}_{{\sf T}^{\rm hol},{t}}+{{\cal I}}\Omega_{\sf T}$. Therefore, ${\sf v}({{\cal I}})$ is in ${\sf v}(\Omega){\cal O}_{{\sf T}^{\rm hol},{t}}+{{\cal I}}$. Since we have ${\sf v}(\Omega)\subset\bar{}{{\cal I}}\subset{{\cal I}}$ we get the result.
@@
If there is an element of $\Theta_{\sf T}(\Omega)$ which is not a zero divisor (in particular if ${\cal O}_{\sf T}$ has no zero divisors and $\Theta\subsetneq\Theta_{\sf T}$) then the following are equivalent:
@@
1. implies 2.: If $L$ is Zariski dense then ¯$\bar{}{{\cal I}}=0$ and this follows from (33). 2. implies 1.: Assume that $L$ is not Zariski dense. The algebraic ideal $\bar{}{{\cal I}}$ has a non-zero element $f$, and $f\Theta_{{\sf T}}\subset\Theta_{{\sf T},L}$. If $\Theta_{{\sf T},L}=\Theta$ then $f\Theta_{{\sf T}}(\Omega)=0$. But by our assumption, $\Theta_{{\sf T}}(\Omega)$ has an element which is not a zero divisor and we get a contradiction. ∎
@@
First, let us define the pairing. For $f,g\in{\cal O}_{\sf T}$ and $\omega=fdg$ we define $\omega(v):=f_{0}v(g-g_{0})$, where $g_{0}=g({t}),\ f_{0}=f({t})$. Since $\Omega_{\sf T}$ as ${\mathfrak{R}}$-module is generated by $fdg$’s, the definition extend to $\Omega^{1}_{\sf T}$. It is well-defined because
@@
If $v\in{{\bf T}}_{t}L$ then we know that ${{\bf T}}_{t}L=\left(\frac{{\mathfrak{m}}_{{\sf T}^{\rm hol},t}}{{{\cal I}}+{\mathfrak{m}}_{{\sf T}^{\rm hol},{t}}^{2}}\right)^{\vee}$, and so we get a map $v:{\cal O}_{{\sf T}^{\rm hol},t}\to{\mathfrak{R}}$ with $v({{\cal I}})=0$. We use the definition of a leaf scheme and we have $\Omega\subset{\cal O}_{{\sf T}^{\rm hol},t}d{{\cal I}}+{{\cal I}}\Omega_{\sf T}$. Applying $v$ to both sides we get $\Omega(v)=0$. Conversely if we have $v\in{{\bf T}}_{t}{\sf T}$ with $\Omega(v)=0$ then we use again the definition of leaf scheme and we have $d{{\cal I}}\subset\Omega\otimes{\cal O}_{{\sf T}^{\rm hol},t}+{{\cal I}}\Omega_{\sf T}$. Applying $v$ to both sides we conclude that $v({{\cal I}})=0$
@@
The evaluation at ${t}$ induces a map $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$ and if $L$ is a general leaf then it is surjective.
@@
Let ${\sf v}\in\Theta_{{\sf T},L}$ and so ${\sf v}({{\cal I}})\subset{{\cal I}}$. This implies that ${\sf v}({{\cal I}})({t})=0$ and so ${\sf v}({t})\in{{\bf T}}_{t}L$. Therefore, the evaluation at ${t}$ induces a map $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$. We have $\Theta\subset\Theta_{{\sf T},L}$, and if $L$ is a general leaf we claim that $\Theta\to{{\bf T}}_{t}L$ is surjective. In order to see this, we first define $\Theta^{\rm hol}:=\{{\sf v}\in\Theta_{{\sf T}^{{\rm hol}},{t}}\ \mid\ {\sf v}(\Omega)=0\}$. Since ${t}$ is a smooth point of ${\sf T}$, we have a holomorphic coordinate system given by the Frobenius theorem and in this coordinate system the surjectivity of $\Theta^{{\rm hol}}\to{{\bf T}}_{t}L$ can be checked easily. The proof finishes with $\Theta\otimes_{{\cal O}_{\sf T}}{{\cal O}_{{\sf T}^{\rm hol},{t}}}=\Theta^{\rm hol}$ which is just a linear algebra.
@@
Note that the kernel of $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$ contains vector fields ${\sf v}$ such that ${\sf v}({t})=0$. In this case ${t}$ is called the singularity of ${\sf v}$. Moreover, for an arbitrary leaf scheme $L$, the map $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$ may not be surjective. For instance, let ${\sf T}:=\mathbb{A}^{2}_{\mathfrak{R}}$ with the coordinates $(x,y)$, $L:xy=0$ and $\Omega={\cal O}_{\sf T}(xdy+ydx)$. It can be easily checked that $L$ is a leaf scheme of ${\cal F}(\Omega)$, $\Theta_{{\sf T},L}=\{xf\frac{\partial}{\partial x}+yg\frac{\partial}{\partial y},\ \ f,g\in{\cal O}_{\sf T}\}$, ${{\bf T}}_{0}L={{\bf T}}_{0}{\sf T}$ and $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$ is the zero map.
@@ 3.6 Foliations of linear differential equations
Let ${V}$ be an ${\mathfrak{R}}$-scheme and ${\sf B}$ be a ${\sf h}\times{\sf h}$ matrix with entries which are global sections of $\Omega^{1}_{V}$. Let also 00 be an ${\mathfrak{R}}$-valued smooth point of ${V}$. Recall that ${\mathfrak{R}}_{\mathbb{Q}}:={\mathfrak{R}}\otimes_{\mathbb{Z}}\mathbb{Q}$ and if ${\mathfrak{R}}\subset\mathbb{C}$ then this is the smallest subring of $\mathbb{C}$ containing both ${\mathfrak{R}}$ and $\mathbb{Q}$. The first fundamental theorem of linear differential equations is:
@@
For any $y_{0}\in{\mathfrak{R}}^{\sf h}$, the linear differential equation $dy={\sf B}y$ has a unique solution $y\in{\cal O}_{{V}_{\mathbb{C}}^{\rm hol},0}^{\sf h}$ with $y(0)=y_{0}$ if and only if ${\sf B}$ satisfies the integrability condition $d{\sf B}={\sf B}\wedge{\sf B}$. Moreover, if ${\sf B}$ is defined over a ring ${\mathfrak{R}}\subset\mathbb{C}$ then $y_{i}$’s are convergent.
@@
The entries of $dx-{\sf B}x$ are differential forms in ${\sf T}:={V}\times_{\mathfrak{R}}O$. If ${\sf B}$ is integrable in the sense of 3.2 then the ${\cal O}_{V}$-module generated by the entries of $dx-{\sf B}x$ is integrable. This follows from
@@
We denote the corresponding foliation in ${\sf T}$ by ${\cal F}(dx-{\sf B}x)$.
@@
All the leaves of ${\cal F}:={\cal F}(dx-{\sf B}x)$ are general and hence smooth.
@@
First note that for a point $(0,y_{0})\in{\sf T}:={V}\times O$, there is a unique solution $y({t})$ of $dy={\sf B}y,\ y(0)=y_{0}$. This gives us the leaf $L$ whose ideal is generated by the entries of $x-y(t)$. The linear part of the entries of $x-y(t)$ is $x$ minus the linear part of $y(t)$ at 00. These are trivially ${\sf h}$ linearly independent functions due to the presence of $x$. ∎
@@
Let $Y\to{V}$ be a family of smooth complex projective varieties over the field ${\mathfrak{k}}$ and let ${V}$ be irreducible, smooth and affine. Around any point of ${V}$ we can find global sections $\omega$ of the ${m}$-th relative de Rham cohomology sheaf of $Y/{V}$ such that $\omega_{i},i=1,2,\ldots,{\sf h}$ at each fiber $H_{\rm dR}^{m}(Y_{t}),\ t\in{V}$ form a basis compatible with the Hodge filtration. If it is necessary we may replace ${V}$ with a Zariski open subset of ${V}$. We write the Gauss-Manin connection of $Y/{V}$ in the basis $\omega=[\omega_{1},\omega_{2},\ldots,\omega_{{\sf h}}]^{{\mathsf{t}}{\mathsf{r}}}$: $\nabla\omega={\sf B}\otimes\omega$, where ${\sf B}$ is ${{\sf h}}\times{{\sf h}}$ matrix with entries which are differential forms in ${V}$. As all our objects $Y\to{V},\ {t},\ \omega_{i}$ etc. use a finite number of coefficients in ${\mathfrak{k}}$, we can take a model of all these over a finitely generated ring ${\mathfrak{R}}$ so that ${\mathfrak{k}}$ is the quotient field of ${\mathfrak{R}}$. For simplicity, we use the same notations for these objects defined over ${\mathfrak{R}}$.
@@
Let ${\mathfrak{R}}\subset\mathbb{C}$. The linear differential equation $dy={\sf B}y$ for an unknown ${{\sf h}}\times 1$ matrix $y$ with entries which are holomorphic functions in $({V},{t})$, has a basis of solutions given by $\int_{\delta_{t}}\omega$, where $\delta_{t}$ ranges in a basis of $H_{m}(Y_{t},\mathbb{Z})$.
This is classical statement in Hodge theory, see for instance or [MV21, Theorem 9.3].

### v2

@@ 3.5 Vector fields tangent to leaf schemes
The formulation of 3.2 with the module $\Theta_{{\sf T},L}$ arises the question how big this module is? For instance, if $\Theta_{{\sf T},L}=0$ then the hypothesis of 3.2 is automatically satisfied and one might doubt this conjecture. From another side, if t{t} is a smooth point of ${\mathcal{F}}(\Omega)$ then $L$ is a general leaf and $\Theta\subset\Theta_{{\sf T},L}$. In this section we gather some statements about $\Theta_{{\sf T},L}$.
[…]
@@
(3.6) $\bar{L}:={\rm Zero}\left(\bar{}{{\mathcal{I}}}\right),\ \ \bar{}{{\mathcal{I}}}:={\mathcal{O}}_{{\sf T}}\cap{{\mathcal{I}}}.$
@@
${\mathcal{O}}_{{\sf T}}\cap{{\mathcal{I}}}=\{0\},$
which means that $\mathcal{I}$}} has not algebraic elements.
[…]
@@
Let $\sf T$} be an $\mathfrak{R}$}-scheme and t{t} be an $\mathfrak{R}$}-valued point of $\sf T$}, not necessarily smooth, and $L$ be a leaf scheme of ${\mathcal{F}}(\Omega)$ through t{t}. We have
(3.7) $\Theta_{{\sf T},L}=\{{\sf v}\in\Theta_{{\sf T}}\ |\ {\sf v}(\Omega)\subset\bar{}{{\mathcal{I}}}\}.$
We may call the right hand side of (3.7) the dual of $\Omega$ along the Zariski closure of $L$. It is a kind of surprising that we do not need to insert ${\sf v}(\bar{}{{\mathcal{I}}})\subset\bar{}{{\mathcal{I}}}$ in (3.7) as it will be clear in the proof.
@@
Proof of $\subset$: If ${\sf v}\in\Theta_{{\sf T},L}$ then ${\sf v}({{\mathcal{I}}})\subset{{\mathcal{I}}}$, and since $\sf v$} is algebraic, that is ${\sf v}({\mathcal{O}}_{\sf T})\subset{\mathcal{O}}_{\sf T}$), we get ${\sf v}(\bar{}{{\mathcal{I}}})\subset\bar{}{{\mathcal{I}}}$. Moreover, by definition of a leaf scheme we have $\Omega={\mathcal{O}}_{\sf T}d{{\mathcal{I}}}$ modulo ${{\mathcal{I}}}\Omega_{\sf T}$ (and after tensoring with holomorphic functions). Taking $\sf v$} from both sides we have ${\sf v}(\Omega)\subset{{\mathcal{I}}}$, and since ${\sf v}(\Omega)\in{\mathcal{O}}_{\sf T}$, we get ${\sf v}(\Omega)\subset\bar{}{{\mathcal{I}}}$.
Proof of $\supset$: Let ${\sf v}\in\Theta_{{\sf T}}$ with ${\sf v}(\Omega)\subset\bar{}{{\mathcal{I}}}$. By definition of a leaf scheme $d{{\mathcal{I}}}$ is in $\Omega\otimes_{{\mathcal{O}}_{\sf T}}{\mathcal{O}}_{{\sf T}^{\rm hol},{t}}+{{\mathcal{I}}}\Omega_{\sf T}$. Therefore, ${\sf v}({{\mathcal{I}}})$ is in ${\sf v}(\Omega){\mathcal{O}}_{{\sf T}^{\rm hol},{t}}+{{\mathcal{I}}}$. Since we have ${\sf v}(\Omega)\subset\bar{}{{\mathcal{I}}}\subset{{\mathcal{I}}}$ we get the result.
@@
If there is an element of $\Theta_{\sf T}(\Omega)$ which is not a zero divisor (in particular if ${\mathcal{O}}_{\sf T}$ has no zero divisors and $\Theta\subsetneq\Theta_{\sf T}$) then the following are equivalent:
@@
1. implies 2.: If $L$ is Zariski dense then $\bar{}{{\mathcal{I}}}=0$ and this follows from (3.7). 2. implies 1.: Assume that $L$ is not Zariski dense. The algebraic ideal $\bar{}{{\mathcal{I}}}$ has a non-zero element $f$, and $f\Theta_{{\sf T}}\subset\Theta_{{\sf T},L}$. If $\Theta_{{\sf T},L}=\Theta$ then $f\Theta_{{\sf T}}(\Omega)=0$. But by our assumption, $\Theta_{{\sf T}}(\Omega)$ has an element which is not a zero divisor and we get a contradiction. ∎
@@
First, let us define the pairing. For $f,g\in{\mathcal{O}}_{\sf T}$ and $\omega=fdg$ we define $\omega(v):=f_{0}v(g-g_{0})$, where $g_{0}=g({t}),\ f_{0}=f({t})$. Since $\Omega_{\sf T}$ as $\mathfrak{R}$}-module is generated by fdgfdg’s, the definition extend to $\Omega^{1}_{\sf T}$. It is well-defined because
@@
If $v\in{{\bf T}}_{t}L$ then we know that ${{\bf T}}_{t}L=\left(\frac{{\mathfrak{m}}_{{\sf T}^{\rm hol},t}}{{{\mathcal{I}}}+{\mathfrak{m}}_{{\sf T}^{\rm hol},{t}}^{2}}\right)^{\vee}$, and so we get a map $v:{\mathcal{O}}_{{\sf T}^{\rm hol},t}\to{\mathfrak{R}}$ with $v({{\mathcal{I}}})=0$. We use the definition of a leaf scheme and we have $\Omega\subset{\mathcal{O}}_{{\sf T}^{\rm hol},t}d{{\mathcal{I}}}+{{\mathcal{I}}}\Omega_{\sf T}$. Applying $v$ to both sides we get $\Omega(v)=0$. Conversely if we have $v\in{{\bf T}}_{t}{\sf T}$ with $\Omega(v)=0$ then we use again the definition of leaf scheme and we have $d{{\mathcal{I}}}\subset\Omega\otimes{\mathcal{O}}_{{\sf T}^{\rm hol},t}+{{\mathcal{I}}}\Omega_{\sf T}$. Applying $v$ to both sides we conclude that $v({{\mathcal{I}}})=0$
@@
The evaluation at t{t} induces a map $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$ and if $L$ is a general leaf then it is surjective.
@@
Let ${\sf v}\in\Theta_{{\sf T},L}$ and so ${\sf v}({{\mathcal{I}}})\subset{{\mathcal{I}}}$. This implies that ${\sf v}({{\mathcal{I}}})({t})=0$ and so ${\sf v}({t})\in{{\bf T}}_{t}L$. Therefore, the evaluation at t{t} induces a map $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$. We have $\Theta\subset\Theta_{{\sf T},L}$, and if $L$ is a general leaf we claim that $\Theta\to{{\bf T}}_{t}L$ is surjective. In order to see this, we first define $\Theta^{\rm hol}:=\{{\sf v}\in\Theta_{{\sf T}^{{\rm hol}},{t}}\ \mid\ {\sf v}(\Omega)=0\}$. Since t{t} is a smooth point of $\sf T$}, we have a holomorphic coordinate system given by the Frobenius theorem and in this coordinate system the surjectivity of $\Theta^{{\rm hol}}\to{{\bf T}}_{t}L$ can be checked easily. The proof finishes with $\Theta\otimes_{{\mathcal{O}}_{\sf T}}{{\mathcal{O}}_{{\sf T}^{\rm hol},{t}}}=\Theta^{\rm hol}$ which is just a linear algebra.
@@
Note that the kernel of $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$ contains vector fields $\sf v$} such that ${\sf v}({t})=0$. In this case t{t} is called the singularity of $\sf v$}. Moreover, for an arbitrary leaf scheme $L$, the map $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$ may not be surjective. For instance, let ${\sf T}:=\mathbb{A}^{2}_{\mathfrak{R}}$ with the coordinates $(x,y)$, $L:xy=0$ and $\Omega={\mathcal{O}}_{\sf T}(xdy+ydx)$. It can be easily checked that $L$ is a leaf scheme of ${\mathcal{F}}(\Omega)$, $\Theta_{{\sf T},L}=\{xf\frac{\partial}{\partial x}+yg\frac{\partial}{\partial y},\ \ f,g\in{\mathcal{O}}_{\sf T}\}$, ${{\bf T}}_{0}L={{\bf T}}_{0}{\sf T}$ and $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$ is the zero map.
@@ 3.6 Foliations of linear differential equations
Let V{V} be an $\mathfrak{R}$}-scheme and $\sf B$} be a ${\sf h}\times{\sf h}$ matrix with entries which are global sections of $\Omega^{1}_{V}$. Let also 0 be an $\mathfrak{R}$}-valued smooth point of V{V}. Recall that $\mathfrak{R}$$_{\mathbb{Q}}$$\mathfrak{R}$$\otimes_{\mathbb{Z}}\mathbb{Q}$ and if $\mathfrak{R}$$\subset\mathbb{C}$ then this is the smallest subring of $\mathbb{C}$ containing both $\mathfrak{R}$} and $\mathbb{Q}$. The first fundamental theorem of linear differential equations is:
@@
For any $y_{0}\in{\mathfrak{R}}^{\sf h}$, the linear differential equation $dy={\sf B}y$ has a unique solution $y\in{\mathcal{O}}_{{V}_{\mathbb{C}}^{\rm hol},0}^{\sf h}$ with $y(0)=y_{0}$ if and only if $\sf B$} satisfies the integrability condition $d{\sf B}={\sf B}\wedge{\sf B}$. Moreover, if $\sf B$} is defined over a ring $\mathfrak{R}$$\subset\mathbb{C}$ then $y_{i}$’s are convergent.
@@
The entries of $dx-{\sf B}x$ are differential forms in ${\sf T}:={V}\times_{\mathfrak{R}}O$. If $\sf B$} is integrable in the sense of 3.2 then the ${\mathcal{O}}_{V}$-module generated by the entries of $dx-{\sf B}x$ is integrable. This follows from
@@
We denote the corresponding foliation in $\sf T$} by ${\mathcal{F}}(dx-{\sf B}x)$.
@@
All the leaves of ${\mathcal{F}}:={\mathcal{F}}(dx-{\sf B}x)$ are general and hence smooth.
@@
First note that for a point $(0,y_{0})\in{\sf T}:={V}\times O$, there is a unique solution y(t)y({t}) of $dy={\sf B}y,\ y(0)=y_{0}$. This gives us the leaf $L$ whose ideal is generated by the entries of xy(t)x-y(t). The linear part of the entries of xy(t)x-y(t) is $x$ minus the linear part of $y(t)$ at 0. These are trivially $\sf h$} linearly independent functions due to the presence of $x$. ∎
@@
Let $Y\to{V}$ be a family of smooth complex projective varieties over the field $\mathfrak{k}$} and let V{V} be irreducible, smooth and affine. Around any point of V{V} we can find global sections $\omega$ of the m{m}-th relative de Rham cohomology sheaf of Y/VY/{V} such that $\omega_{i},i=1,2,\ldots,{\sf h}$ at each fiber $H_{\rm dR}^{m}(Y_{t}),\ t\in{V}$ form a basis compatible with the Hodge filtration. If it is necessary we may replace V{V} with a Zariski open subset of V{V}. We write the Gauss-Manin connection of Y/VY/{V} in the basis $\omega=[\omega_{1},\omega_{2},\ldots,\omega_{{\sf h}}]^{{\mathsf{t}}{\mathsf{r}}}$: $\nabla\omega={\sf B}\otimes\omega$, where $\sf B$} is ${{\sf h}}\times{{\sf h}}$ matrix with entries which are differential forms in V{V}. As all our objects $Y\to{V},\ {t},\ \omega_{i}$ etc. use a finite number of coefficients in $\mathfrak{k}$}, we can take a model of all these over a finitely generated ring $\mathfrak{R}$} so that $\mathfrak{k}$} is the quotient field of $\mathfrak{R}$}. For simplicity, we use the same notations for these objects defined over $\mathfrak{R}$}.
@@
Let $\mathfrak{R}$$\subset\mathbb{C}$. The linear differential equation $dy={\sf B}y$ for an unknown ${{\sf h}}\times 1$ matrix $y$ with entries which are holomorphic functions in (V,t)({V},{t}), has a basis of solutions given by $\int_{\delta_{t}}\omega$, where $\delta_{t}$ ranges in a basis of $H_{m}(Y_{t},\mathbb{Z})$.
This is classical statement in Hodge theory, see for instance or [ho13-Roberto, Theorem 9.3].